# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

### v4

@@ II Preliminaries
For a thorough introduction to Tweedie’s formula, we refer the reader to [efron_tweedies]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms.
Theorem II.9 ([gribonval_should] and [gribonval_reconciling, Corollary 1]).
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.333i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE estimator under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
  1. 1.
@@
    where $f_{Z}=-\log p_{Z}$.
  4. 4.
    $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$.
@@ III Main Results
In this section, we present our theoretical contributions. We begin by introducing a new structural representation of the implicit regularizer underlying the MMSE denoiser. This characterization provides greater interpretability and serves as the foundation for deriving novel nonasymptotic convergence guarantees.
III-A A New Characterization of the MMSE Regularizer.
It is known (Theorem II.9) that the MMSE denoiser is the proximal map of an infinitely differentiable penalty. Yet beyond this, the structure of the associated penalty $\phi_{\mathrm{MMSE}}$ remains unclear. The only known explicit expression of $\phi_{\mathrm{MMSE}}$ (28) is cumbersome and offers little analytical insight. In this section, we show that $\phi_{\mathrm{MMSE}}$ admits an upper Moreau envelope representation in terms of $f_{Z}$, yielding useful insights, namely $1$-weak convexity, which we will leverage to prove stronger convergence results for the PGD using an MMSE denoiser.
@@
Pick $x\in\mathbb{R}^{n}$. By Theorem II.8,
@@
$=\sigma^{2}\nabla f_{Z}(x),$ (31)
which gives us
$\sigma^{2}\nabla f_{Z}(x)=\nabla M_{1}\phi_{\mathrm{MMSE}}(x),$ (32)
when combined with Theorem II.7. Integrating gives the result. ∎
@@
The regularizer implicit in the MMSE denoiser can be written as
@@
For the reverse inequality, take some $x$ in the image of $\psi_{\sigma}$. There exists some $z\in\mathbb{R}^{n}$ so that $x=\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$. The infimum defining $M_{1}\phi_{\mathrm{MMSE}}$ is achieved at $x$ (see (18)), so that
@@
$\leq\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}}.$ (37)
We illustrate this result in Fig. 1 and Fig. 2 under a variety of mixture of Gaussian and mixture of Laplacian priors. An overview of the experiment is provided in Section IV-A.
Figure 1: Calculated and learned regularizers for an MMSE denoiser under mixture-of-Gaussian priors with unit Gaussian noise. The calculated regularizer is derived via Theorem III.2.
Figure 2: Calculated and learned regularizers for an MMSE denoiser under mixture-of-Laplacian priors with unit Gaussian noise. The calculated regularizer is derived via Theorem III.2.
Corollary III.3.
The regularizer $\phi_{\mathrm{MMSE}}$ implicit in the MMSE denoiser is $1$-weakly convex on the image of $\psi_{\sigma}$.
@@
From Theorem III.2,
$\phi_{\mathrm{MMSE}}$ $=\sigma^{2}M^{\sigma^{2}}f_{Z}-\sigma^{2}C_{x_{0}}$
$=M^{1}(\sigma^{2}(f_{Z}-C_{x_{0}}))=M^{1}M_{1}\phi_{\mathrm{MMSE}},$ (38)
by Section III-A. This holds only if $\phi_{\mathrm{MMSE}}$ is 1-weakly convex by Section II. ∎
We now leverage this 1-weak convexity to derive nonasymptotic guarantees for PnP-PGD.
III-B Proximal Gradient Descent Convergence Results.
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient and subgradient (see [rockafellar_variational, Chapter 8]) may not exist and are generalized to the Fréchet subgradient.
Definition III.4 (Fréchet Subgradient).
For a function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and a point $x$ where $F$ is finite, the Fréchet Subgradient $\widehat{\partial}F(x)$ is given by
$\widehat{\partial}F(x)\coloneqq\mathopen{}\mathclose{{\left\{v\in\mathbb{R}^{n}\;\middle|\;\liminf_{y\to x}\frac{f(x)-f(y)-\langle v,x-y\rangle}{\|x-y\|}}}\right\}.$ (39)
Informally, $\widehat{\partial}F(x)$ is the set of all planes that locally support $F$ at $x$. $x^{\star}$ is a stationary point for $F$ if and only if $0\in\widehat{\partial}F(x^{\star})$, and when $F$ is differentiable, $\widehat{\partial}F(x)=\{\nabla F(x)\}$. This motivates measuring progress by approximate stationarity. In the nonsmooth, non-convex setting, the natural goal is to find an $\varepsilon$-stationary point, i.e. an $x$ satisfying
$\operatorname{dist}(0,\widehat{\partial}F(x))\leq\varepsilon,$ (40)
@@
In non-convex, nonsmooth problems, the stationary residual (41) may oscillate, even while it trends downward. This can depend on the forward operator as well as the data. To account for this behavior, nonasymptotic convergence results are typically stated in terms of the best iterate so far. See for example [attouch_convergence]. It is standard to state that an algorithm reaches an $\varepsilon$-stationary point in the first $k$ iterations if
$\min_{1\leq j\leq k}\operatorname{dist}(0,\widehat{\partial}F(x_{j}))\leq\varepsilon.$ (42)
For our investigation, we prove the following intermediary result, which is a simple improvement over [bohm_variable, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser. For our immediate application, $\widehat{\partial}F(x)$ reduces to $\{\nabla F(x)\}$, but we state the result in the most general form because it provides a unified notion of stationarity that remains valid in nonsmooth and nonconvex settings.
Lemma III.5.
Let $\{x_{k}\}$ be iterates of proximal gradient descent defined in (5), where $g$ is proper, lower-semicontinuous, $\rho$-weakly convex, and $\nabla f$ is $L$-Lipschitz continuous. Suppose $\gamma\in\mathopen{}\mathclose{{\left(0,\tfrac{2}{L+\rho}}}\right)$. Then, for all $k\geq 1$,
$\min_{2\leq j\leq k+1}\operatorname{dist}(0,\widehat{\partial}F(x_{j}))\leq k^{-1/2}C\sqrt{F(x_{1})-F^{*}},$ (43)
where $F$ is as in (3),
@@
and
$C=\tfrac{\tfrac{1}{\gamma}+L}{\sqrt{\tfrac{1}{\gamma}-\tfrac{L+\rho}{2}}}.$ (45)
In other words, there is an $\epsilon$-stationary point in the first $k+1$ iterates with $\epsilon\sim\mathcal{O}(1/\sqrt{k})$.
Proof.
We first note that $\widetilde{g}(x)\coloneqq g(x)+\tfrac{\rho}{2}\|x\|^{2}$ is convex, so
@@
for any $s\in\widehat{\partial}\widetilde{g}(x)$. For iteration $x_{k+1}$, let $s_{k+1}\in\widehat{\partial}g(x_{k+1})$, so that $s_{k+1}+\rho x_{k+1}\in\widehat{\partial}\widetilde{g}(x_{k+1})$. We then have
$g(x_{k})+\tfrac{\rho}{2}\|x_{k}\|^{2}$ $\geq g(x_{k+1})+\tfrac{\rho}{2}\|x_{k+1}\|^{2}$
$+\langle s_{k+1}+\rho x_{k+1},x_{k}-x_{k+1}\rangle,$ (47)
using Eq. 46 with $y=x_{k}$, $x=x_{k+1}$, and $s=s_{k+1}+\rho x_{k+1}$. We will write this as
@@
$-\tfrac{\rho}{2}\|x_{k}-x_{k+1}\|^{2}.$ (48)
$\nabla f$ is $L$-Lipschitz, so
@@
$+\tfrac{L}{2}\|x_{k+1}-x_{k}\|^{2}.$ (49)
Setting $s_{k+1}\coloneq\tfrac{1}{\gamma}(x_{k}-x_{k+1})-\nabla f(x_{k})\in\widehat{\partial}g(x_{k+1})$ and adding (III-B) and (III-B),
[…]
Next, let $w_{k+1}\coloneqq\tfrac{1}{\gamma}(x_{k+1}-x_{k})+\nabla f(x_{k+1})-\nabla f(x_{k})$, an element of $\widehat{\partial}(f+g)(x_{k+1})$,
$\|w_{k+1}\|^{2}$ $\leq\mathopen{}\mathclose{{\left(\tfrac{1}{\gamma}+L}}\right)^{2}\|x_{k+1}-x_{k}\|^{2}$
$\leq C^{2}\mathopen{}\mathclose{{\left(F(x_{k})-F(x_{k+1})}}\right),$ (51)
by (50), where $C=\tfrac{\tfrac{1}{\gamma}+L}{\sqrt{\tfrac{1}{\gamma}-\tfrac{L+\rho}{2}}}$. Next, we have a simple telescoping sum,
$\min_{1\leq j\leq k}\|w_{j+1}\|^{2}$ $\leq\tfrac{1}{k}\sum_{j=1}^{k+1}\|w_{j}\|^{2}$
$\leq\tfrac{1}{k}C^{2}(F(x_{1})-F(x_{k+1}))$
$\leq C^{2}(F(x_{1})-F^{*}),$ (52)
using the definition in (44). Taking a square root and recalling that $\operatorname{dist}(0,\widehat{\partial}F(x_{j+1}))\leq\|w_{j+1}\|$ gives us the result. ∎
Remark III.6.
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [xu_provable] showed that PnP-MMSE converges for any $\gamma\leq 1/L$. Though the MMSE estimator can be written as $\operatorname{prox}_{\phi{\mathrm{MMSE}}}$, it is impossible to evaluate $\operatorname{prox}_{\gamma\phi{\mathrm{MMSE}}}$ for $\gamma\neq 1$. The denoiser fixes the effective step size. Consequently, any change in the step size is really just a rescaling of the data-fidelity term. For this reason, we assume an implicit step size of $1$ and require the Lipschitz constant of the data-fidelity term to be sufficiently small.
Theorem III.7.
Consider the PnP iteration (6) using $D_{\sigma}\coloneqq\psi_{\sigma}$, the MMSE denoiser with implicit step size $\gamma=1$. If $\nabla f$ is $L$-Lipschitz continuous with $L<1$, there exists an $\mathcal{O}(1/\sqrt{k})$-stationary point among $\{x_{j}\}_{j=1}^{k+1}$.
Proof.
Apply Section III-B with $\rho=1$, since $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. Note that the iterates $x_{j}$ lie in the image of $\psi_{\sigma}$ for $j\geq 2$, so Theorem III.2 applies to them. ∎
IV Experiments
IV-A Illustration of Theorem III.2.
Figure 3: MMSE denoiser, convex potential, and implicit MMSE regularizer for a Laplacian prior $p_{X}(x)=e^{-|x|}$, plotted against reference curves. LPN denotes the learned proximal network. This explicates what is shown in [fang_whats, Figure 2], but plotted on a wider axis to show the “curling down” behavior of LPNs.
We consider the denoising problem (2) under mixture-of-Gaussian and mixture-of-Laplacian priors with unit Gaussian noise. We train a gradient-output softplus input convex neural network (ICNN) to approximate the MMSE denoiser and calculate its implicit prior via [gribonval_characterization, Theorem 3]. We compare this with the implicit regularizer calculated using the explicit upper Moreau envelope formula from Theorem III.2. Figures 1 and 2 compare the calculated $\phi_{\mathrm{MMSE}}$ and the learned estimate, along with reference curves $f_{X}\coloneqq-\log p_{X}$ and $f_{Z}\coloneqq-\log p_{Z}$. The learned and calculated regularizers are nearly identical where the samples are most concentrated.
We observe that the learned implicit regularizer “curls downward” outside the central region of the data. This behavior is intrinsic to ReLU-based neural networks (and their softplus approximations). It is well known that feedforward networks with continuous piecewise-affine activations yield continuous piecewise-affine mappings [raghu_expressive, montufar_number]. The ReLU is piecewise-affine and is the uniform limit of the softplus family. Thus, an ICNN with softplus activations is a smooth approximation of a ReLU-ICNN and is therefore nearly piecewise-affine.
Outside the support of the training data, the learned convex potential becomes approximately affine, and its gradient (the denoiser) saturates. To solve for $\phi_{\mathrm{MMSE}}(x)$, we must invert the learned denoiser numerically. This inversion is impossible where the denoiser saturates, so we get large values for $\psi_{\theta}^{-1}$ outside the central region. We solve for $\phi_{\mathrm{MMSE}}(x)$ using  [gribonval_characterization, Eq. (8)], so the quadratic $g_{\theta}(\psi_{\theta}^{-1}(x))$ dominates, causing the resulting estimate of the regularizer to curl downward to -$\infty$. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [fang_whats, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
IV-B Gaussian Blurring.

### v6

@@ II Preliminaries
For a thorough introduction to Tweedie’s formula, we refer the reader to [22]. Note that Tweedie’s Formula makes no assumptions on the prior density $p_{X}$, which may even include discrete atoms. We recall the following result on MMSE denoisers due to Gribonval et al. [29, 31].
Theorem II.8 (Main result from [31] and [29, Corollary 1]).
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.222i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE denoiser under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
  1. (i)
@@
  1. where $f_{Z}=-\log p_{Z}$.
  2. (iv)
    $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$ on $\psi_{\sigma}(\mathbb{R}^{n})$.
@@ III Main Results
In this section, we present our theoretical contributions. We begin by introducing a new structural representation of the implicit regularizer underlying the MMSE denoiser. This characterization serves as the foundation for deriving novel nonasymptotic convergence guarantees.
III-A New Characterization of the MMSE Regularizer
It is known (Theorem II.8) that the MMSE denoiser is the proximal map of an infinitely differentiable penalty. Yet beyond this, the structure of the associated penalty $\phi_{\mathrm{MMSE}}$ remains unclear. The only known explicit expression of $\phi_{\mathrm{MMSE}}$ (24) is cumbersome and offers little analytical insight. In this section, we show that $\phi_{\mathrm{MMSE}}$ admits an upper Moreau envelope representation in terms of $f_{Z}$, yielding useful insights, namely $1$-weak convexity, which we will leverage to prove stronger convergence results for the PGD using an MMSE denoiser.
@@ III-A New Characterization of the MMSE Regularizer
Pick $x\in\mathbb{R}^{n}$. By Theorem II.7 and Theorem II.8,
@@
$=\sigma^{2}\nabla f_{Z}(x).$ (27)
Combined with Theorem II.6 yields
$\sigma^{2}\nabla f_{Z}(x)=\nabla M_{1}\phi_{\mathrm{MMSE}}(x).$ (28)
Integrating gives the result. ∎
@@
The implicit regularizer in the MMSE denoiser can be written as
@@
For the reverse inequality, take some $x\in\psi_{\sigma}(\mathbb{R}^{n})$. There exists some $z\in\mathbb{R}^{n}$ so that $x=\psi_{\sigma}(z)=\operatorname{prox}_{\phi_{\mathrm{MMSE}}}(z)$. The infimum defining $M_{1}\phi_{\mathrm{MMSE}}$ is achieved at $x$ (cf. (14)), so that
@@
$\leq\sigma^{2}M^{\sigma^{2}}f_{Z}(x)-\sigma^{2}C_{x_{0}},$ (33)
which completes the proof. ∎
We illustrate this result in Fig. 1 and Fig. 2 under a variety of mixture-of-Gaussian and mixture-of-Laplacian priors. An overview of the experimental details is provided in Section IV-A.
Figure 1: Top: Calculated implicit regularizer in the MMSE denoiser ($\phi_{\mathrm{MMSE}}$), negative log-marginal $f_{Z}$, and negative log-prior $f_{X}$ for three mixture-of-Gaussian priors under unit Gaussian noise. Bottom: The learned regularizer $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$, along with the same reference curves. Details in Section IV-A.
Figure 2: Top: Calculated implicit regularizer in the MMSE denoiser ($\phi_{\mathrm{MMSE}}$), negative log-marginal $f_{Z}$, and negative log-prior $f_{X}$ for three mixture-of-Laplacian priors under unit Gaussian noise. Bottom: The learned regularizer $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$, along with the same reference curves. Details in Section IV-A.
III-B Proximal Gradient Descent Convergence Results
By Theorem III.2, $\phi_{\mathrm{MMSE}}$ equals $\sigma^{2}M^{\sigma^{2}}f_{Z}$ on $\psi_{\sigma}(\mathbb{R}^{n})$. By Section II, $M^{\sigma^{2}}f_{Z}$ is $\tfrac{1}{\sigma^{2}}$-weakly convex. Hence, $\sigma^{2}M^{\sigma^{2}}f_{Z}$ is $1$-weakly convex. Therefore, $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex on $\psi_{\sigma}(\mathbb{R}^{n})$. We leverage this to derive nonasymptotic convergence guarantees for PnP-PGD with MMSE denoisers.
Consider the iteration
[…]
where $\psi_{\sigma}$ is the MMSE denoiser under Gaussian noise of variance $\sigma^{2}$ and $f$ is an $L$-smooth data-fidelity term. The step size is fixed because it is implicit in the MMSE denoiser. The associated objective for these iterates is
[…]
where $\phi_{\mathrm{MMSE}}$ is the implicit regularizer in the MMSE estimator, discussed in Section III-A.
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural as small gradient norms imply closeness to a stationary point. For nonconvex problems, $\|\nabla F(x_{k})\|$ may oscillate even as it trends downward, so nonasymptotic convergence is typically stated in terms of the best iterate up until the current iteration [4]. Since $f$ is smooth and $\phi_{\mathrm{MMSE}}$ is $C^{\infty}$ on its domain (Theorem II.8), the objective $F$ is $C^{\infty}$ on its domain. We measure convergence via $\min_{i\leq K}\|\nabla F(x_{i})\|$ at each iteration $K$.
Theorem III.3 (Convergence of Eq. 34).
Let $f$ be proper, bounded from below and $L$-smooth with $L<1$. Then for $F$ in Eq. 35, the iterates of (34) satisfy the following:
  1. (i)
    The sequence $(F(x_{k}))$ is non-increasing and converges.
  2. (ii)
    The sequence has finite length, i.e., $\sum_{k=0}^{\infty}\|x_{k+1}-x_{k}\|^{2}<\infty$, and $\|x_{k+1}-x_{k}\|$ converges to 0 at rate
    $\min_{i\leq K}\|x_{i+1}-x_{i}\|=\mathcal{O}\mathopen{}\mathclose{{\left(\tfrac{1}{\sqrt{K}}}}\right).$ (36)
  3. (iii)
    All cluster points of $x_{k}$ are stationary points of $F$.
@@ III-B Proximal Gradient Descent Convergence Results
We have that $\sigma^{2}M^{\sigma^{2}}f_{Z}-\sigma^{2}C_{x_{0}}$ is a globally-defined 11-weakly convex function that agrees with $\phi_{\mathrm{MMSE}}$ on $\psi_{\sigma}(\mathbb{R}^{n})$. Next, each iterate of (34) clearly satisfies $x_{k}\in\psi_{\sigma}(\mathbb{R}^{n})$ for $k\geq 1$. The application of [34, Theorem 1] with $\tau=1$, $M=1$, and $L_{f}<1$ yields the result. ∎
Corollary III.4.
Under the assumptions of Theorem III.3,
$\min_{i\leq K}\|\nabla F(x_{i})\|=\mathcal{O}\mathopen{}\mathclose{{\left(\tfrac{1}{\sqrt{K}}}}\right).$ (37)
Proof.
From Theorem II.8 and Theorem III.2,
$x_{k+1}=\operatorname*{arg\,min}_{z}G_{k}(z),$ (38)
@@
$G_{k}(z)$ $\coloneqq\sigma^{2}(M^{\sigma^{2}}f_{Z}(z)-C_{x_{0}})$
$\quad+\tfrac{1}{2}\|z-(x_{k}-\nabla f(x_{k}))\|^{2}.$ (39)
By Section II, $\sigma^{2}M^{\sigma^{2}}f_{Z}$ is $1$-weakly convex. Hence, $G_{k}$ is convex and we can conclude that its minimizer satisfies first-order optimality, i.e., $\nabla G_{k}(x_{k+1})=0$. We can write this as $\nabla\phi_{\mathrm{MMSE}}(x_{k+1})+x_{k+1}-(x_{k}-\nabla f(x_{k}))=0,$ because $\phi_{\mathrm{MMSE}}(z)$ and $\sigma^{2}M^{\sigma^{2}}f_{Z}(z)-\sigma^{2}C_{x_{0}}$ agree when $z=x_{k+1}$. Therefore,
@@
From this, we have that
$\|\nabla F(x_{k+1})\|$ $\leq\|x_{k}-x_{k+1}\|+\|\nabla f(x_{k+1})-\nabla f(x_{k})\|$
$\leq 2\|x_{k+1}-x_{k}\|.$ (41)
The result then follows from Theorem III.3 (ii). ∎
Remark III.5.
We note that it may seem that the requirement of $L<1$ in Theorem III.3 is restrictive, it turns out that any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment (Section IV-D), the forward operator is a discretized Radon transform, which has an $L$-smooth $f$, where $L\approx 140$. Thus, we simply scale the data-fidelity term to obtain empirical results that verify our theoretical convergence rate.
Remark III.6.
The asymptotic convergence of PGD using an MMSE denoiser has previously been established for a range of step sizes $\gamma>0$. In particular, [65] showed that PnP-MMSE converges for any $\gamma\leq 1/L$. Though the MMSE denoiser can be written as $\operatorname{prox}_{\phi_{\mathrm{MMSE}}}$, we note that $\operatorname{prox}_{\gamma\phi{\mathrm{MMSE}}}$ cannot be evaluated for any $\gamma\neq 1$. Consequently, any change in the step size can be realized by a rescaling of the data-fidelity term. For this reason, we assume an implicit step size of $1$ and scale the Lipschitz constant of the data-fidelity term accordingly.
IV Experiments
In this section, we numerically verify the theoretical results presented in Section III. The code to reproduce our experiments is publicly available on GitHub. [^3] https://github.com/sparsity-group/pnpmmse All training and testing was done on an NVIDIA RTX PRO 6000 Blackwell Max-Q Workstation Edition GPU using the Ubuntu 24.04 Windows Subsystem for Linux.
Central to our experiments is the task of estimating the MMSE denoiser of a given distribution under Gaussian noise. Following the work of [24], we choose to model the MMSE denoiser as the gradient of a softplus deep neural network (DNN). We found that this architecture performed better than rectified linear unit (ReLU) alternatives. We also incorporate skip connections at each layer, as they have been shown to aid in image denoising tasks [43]. Finally, by Tweedie’s formula (Theorem II.7), the true MMSE denoiser can be written as the gradient of $\tfrac{1}{2}\|\cdot\|^{2}+\log p_{Z}$, where $p_{Z}$ is the marginal distribution under Gaussian noise. For this reason, our choice of modeling the MMSE denoiser as the gradient of a DNN is well-motivated.
TABLE II: DNN architectural details for inverse problems experiments (Section IV-C and Section IV-D). $H$ denotes the hidden dimension and $L$ denotes the number of layers.
Gaussian Deblurring Computed Tomography
Input size $28\times 28$ $512\times 512$
Dimensions $H\times 28\times 28,\ H\times 14\times 14,\ H\times 14\times 14,$ $H\times 512\times 512,\ H\times 256\times 256,\ H\times 256\times 256,\ H\times 128\times 128$
$H\times 7\times 7,\ H\cdot 7^{2}\times 1\times 1,\ 64\times 1\times 1$ $H\times 128\times 128,\ H\times 64\times 64,\ H\times 64\times 64,\ H\times 32\times 32,$
$H\times 16\times 16,\ 64\times 1\times 1,\ 64\times 1\times 1$
$L$ 66 $11$
$H$ $64$ $256$
Direct connections ($W_{k}$) 4 Conv2d ($3\times 3$ kernel), 2 Linear 9 Conv2d ($3\times 3$), 1 Conv2d ($16\times 16$ kernel), 1 Linear
Skip connections ($A_{k}$) 3 Conv2d, 1 Linear 8 Conv2d ($3\times 3$), 1 Conv2d ($16\times 16$ kernel)
Figure 3: Best-iterate gradient norm and iterate difference over $100$ iterations of PnP-PGD with an MMSE denoiser for the MNIST Gaussian deblurring experiment in Section IV-C, plotted on a log-log scale. Both quantities decay sublinearly, consistent with a $\mathcal{O}(1/\sqrt{K})$ rate.
IV-A Numerical Illustration of Theorem III.2
Figure 4: PnP-PGD reconstruction results on three selections from the MNIST dataset under Gaussian blur (Section IV-C). From left to right: clean reference, blurred and noisy input, and reconstruction after 100 iterations. PSNR gains are roughly 13131414 dB are achieved across all three examples.
Figure 5: Best-iterate gradient norm and iterate difference over $1000$ iterations of PnP-PGD with an MMSE denoiser for the MayoCT Computed Tomography experiment in Section IV-D, plotted on a log-log scale. Both quantities decay sublinearly, consistent with a $\mathcal{O}(1/\sqrt{K})$ rate.
We first consider the denoising problem Eq. 2 with simple 1-dimensional mixture-of-Gaussian and mixture-of-Laplacian priors, denoted by $p_{X}$, with additive zero-mean, unit-variance Gaussian noise $\epsilon$. In each case, $p_{Z}$ denotes the marginal distribution of $z$. We now describe the DNN architecture used to learn the MMSE denoiser. Let
$h_{0}$ $\coloneqq\sigma(W_{0}x),\quad W_{0}\in\mathbb{R}^{H},$
$h_{k+1}$ $\coloneqq\sigma(W_{k+1}h_{k}+a_{k}x+b_{k}),\quad k=0,\ldots,L-1,$ (42)
and $g_{\theta}(x)\coloneqq w_{\mathrm{out}}^{\mathsf{T}}h_{3}+a_{\mathrm{out}}x+b_{\mathrm{out}}$, where $W_{k}\in\mathbb{R}^{H\times H}$ are the weights, $a_{k},b_{k}\in\mathbb{R}^{H}$ are the skip connections and biases, and $w_{\mathrm{out}},a_{\mathrm{out}},b_{\mathrm{out}}\in\mathbb{R}$ are the output weights and biases.
We use a softplus activation function
@@ IV Experiments
with $\beta=10$. There are $L=4$ hidden layers, with a hidden dimension of $H=50$. In particular, we model our MMSE denoiser as the gradient of the DNN, i.e., $\psi_{\theta}\coloneqq\nabla g_{\theta}$. To learn the denoiser, we optimize the objective
@@
with a batch size $B=2000$, $x_{i}\sim p_{X}$, and $z_{i}=x_{i}+\epsilon$, $\epsilon\sim\mathcal{N}(0,1)$. We consider the cases where $p_{X}$ is either a mixture-of-Gaussian or mixture-of-Laplacian distribution. We trained the DNNs for $500$ epochs with a learning rate $10^{-3}$ followed by $500$ epochs with a learning rate of $10^{-4}$.
As discussed in Section III, the MMSE denoiser $\psi_{\sigma}$ can be written as the $\operatorname{prox}_{\phi_{\mathrm{MMSE}}}$, where $\phi_{\mathrm{MMSE}}$ is the implicit regularizer. To approximate this, we take advantage of [30, Eq. (8)] and write
@@
where we compute $\psi_{\theta}^{-1}$ numerically via least-squares. When $\psi_{\theta}$ is the bona fide MMSE denoiser, $\phi^{\mathrm{learned}}_{\mathrm{MMSE}}$ is the bona fide implicit regularizer [30].
Figure 6: PnP-PGD reconstruction result on three axial slices from the MayoCT dataset. From left to right: clean reference, noisy sinogram, filtered back projection of the noisy sinogram, and PnP-PGD output using an MMSE denoiser with early stopping. PSNR and SSIM are reported for the noisy and denoised images. Reconstructions achieve PSNR gains of roughly 15–20 dB over the noisy input, with SSIM improvements around $0.65$.
Finally, we can also exactly calculate the true MMSE denoiser, $\psi_{\mathrm{MMSE}}$, via Tweedie’s formula (Theorem II.7). From Eq. 14 and Section III-A, we have that
@@
which allows us to solve for $\phi_{\mathrm{MMSE}}$ explicitly.
In Figs. 1 and 2 we compare the learned vs. the calculated implicit regularizers for three mixture-of-Gaussian and three mixture-of-Laplacian priors, along with the reference curve $f_{X}\coloneq-\log p_{X}$. Since $\phi_{\mathrm{MMSE}}$ is only determined up to the additive constant $C_{x_{0}}$ in Section III-A, we shift the curves vertically for visual comparison.
IV-B DNN Architecture for Inverse Problems
In this subsection, we now detail the DNN architecture for the inverse problems experiments that appear in Section IV-C (Gaussian deblurring) and Section IV-D (computed tomography). In both cases, we define $h_{1}\coloneqq\sigma(W_{1}x)$, and let
$h_{k}\coloneqq\sigma(W_{k}h_{k-1}+A_{k-1}\mathcal{I}_{k}(x)+b_{k}),$ (47)
where $k=2,\ldots,(L-1)$ and $\mathcal{I}_{k}(\cdot)$ denotes bilinear interpolation of $x$ so that the dimension match for matrix-vector multiplication. We let the scalar output be $g_{\theta}(x)$, defined as
$w_{\mathrm{out}}^{\mathsf{T}}\sigma(W_{L}\operatorname{vec}(h_{L-1})+A_{L}\operatorname{vec}(\mathcal{I}_{L}(x))+b_{L}).$ (48)
Each layer is either a convolutional or fully-connected layer. There is a skip-connection at every layer except the first and last. The DNNs use a softplus activation Eq. 43 with $\beta=10$. We provide the exact architectural details in Table II. The DNN is trained with the mean-squared error loss Eq. 44 with $z_{i}=x_{i}+\epsilon$ for zero-mean Gaussian noise $\epsilon$ with variance $\sigma^{2}=0.04$.
IV-C Gaussian Deblurring